# Kotko,I.,${}^{1}$ Banerjee,S.,${}^{2,3}$ Belczynski,K.,${}^{1}$

### v1

@@ Abstract
The two systems that have been confirmed as dormant (no X-ray emission detected) black hole (BH) - low-mass star binaries (Gaia BH1 and Gaia BH2) in the latest Gaia mission data release (DR3) are very intriguing in the context of their evolution. Both systems consist of $\sim 9\,\mathrm{M_{\odot}}$ BH and $\sim 1\,\mathrm{M_{\odot}}$ star orbiting each other on a wide, eccentric orbit ($e\sim 0.5$). We argue that the formation of Gaia BH-like systems through the isolated binary evolution (IBE) channel under standard common envelope assumptions and from dynamical interactions in young massive and open clusters is equally probable, and the formation rate of those binaries is on order of $10^{-7}\,\mathrm{M_{\odot}}^{-1}$ in both channels.
What plays an important role in formation of Gaia BH-like systems in the case of IBE channel is the mutual position of the natal kick velocity vector and the binary angular momentum vector. We find that the natal kicks of median magnitude $\sim 40$ km/s are preferred for the formation of Gaia BH1-like binaries. $\sim 94\%$ of those binaries is formed with BH spin misaligned to orbital axis by less than $40^{\circ}$. Gaia BH2-like binaries form if the low velocity natal kick (of median magnitude $\sim 20$ km/s) is directed within $15^{\circ}$ about the orbital plane.
We follow the subsequent evolution of the binaries once formed as Gaia BH1 and Gaia BH2 systems to investigate their connection with the low-mass X-ray binary population.
@@ 1 Introduction
The dormant black hole (BH) binaries are difficult to observe as they do not emit X-ray radiation and can be discovered only by astrometric and/or spectroscopic analysis. Although such systems were predicted to exist already in the 1960s (Guseinov & Zel’dovich, 1966) only recently the new generation of the astronomical instruments and new data analysis methods have allowed to identify the first candidates for such objects. To our knowledge the first strong candidate for a binary system with a heavy companion unseen in electromagnetic spectrum was found in the globular cluster NGC 3201 Giesers et al. (2018). Within last two years another four systems have been discovered: (i) binary HD 130298 where the minimum mass of an unseen companion has been estimated to be $7.7\pm 1.5$ $\mathrm{M_{\odot}}$ (Mahy et al., 2022), (ii) VFTS 243 found in Large Magellanic Cloud which is a binary of massive star and most probably black hole companion with mass $M=10.1\pm 2.0$ $\mathrm{M_{\odot}}$ (Shenar et al., 2022) and (iii) two systems which were identified in the most recent Gaia catalogue DR3 for which there is the strong evidence that they host a non-accreting BH and a low-mass companion star (El-Badry et al., 2023a, b; Tanikawa et al., 2023b). In our studies we focus only on the last two systems and following El-Badry et al. (2023b) we refer to those binaries as Gaia BH1 and Gaia BH2.
@@
Once Gaia BH1 and Gaia BH2 had been detected, the question arose about the origin of those systems. Already several studies have been made which point towards evolution through the dynamical interactions in young stellar clusters as the most probable evolution scenario to create BH-low-mass star system on a moderately wide and eccentric orbit (El-Badry et al., 2023a; Tanikawa et al., 2023a; Rastello et al., 2023; Niccolò Di Carlo et al., 2023). However, as we want to point out in this work, the isolated binary evolution (IBE) channel may contribute to the total population of dormant BH binaries at a rate comparable to the dynamical interactions channel. The important factor that impacts the number of the binaries of interest that may be born through IBE channel is the magnitude and geometry of the natal kick imparted to the newly formed compact object after supernova (SN) explosion. One can infer the information about the natal kick from the peculiar velocity [^1] The peculiar velocity corresponds to systemic velocity in the model, as is explained in Sec. 2.1. of the system and the system orbital parameters.
@@
There are two components that change the systemic velocity of a binary after SN event in the model: (i) the Blaauw kick that impacts the velocity of the binary centre of gravity due to symmetric mass ejection during SN (Blaauw, 1961), (ii) the natal kick that is received by the compact object due to asymmetries of mass ejecta and/or neutrino emission during SN (Chugai, 1984; Janka, 2013, 2017, e.g.). The observations of proper motions of pulsars and X-ray binaries give constrains on the natal kicks received by neutron stars (NSs) or BHs upon their formation. The Maxwellian distribution with dispersion velocity $\sigma=265$ km/s derived by Hobbs et al. (2005) from observations of pulsars is widely used in the binary population synthesis simulations. This distribution is applied also to BHs but with additional assumption that the kick velocity is reduced by the velocity of the matter that falls back pulled by the gravitation field of BH forming during SN. However, recent observations indicate that this approach may underestimate the number of lower velocity systems and overestimate the number of the systems in higher velocity range and that it should be revisited (Atri et al., 2019; O’Doherty et al., 2023; Zhao et al., 2023). The magnitude of the natal kicks is only one part of a puzzle, the other is the kick velocity vector orientation. In the codes based on BSE code from Hurley et al. (2000) the direction of natal kicks is drawn from isotropic spherical distribution. The question is whether the specific direction of natal kicks that BHs receive in SN explosion could lead to the formation of Gaia BH1 and Gaia BH2 (under the standard assumptions of isolated binary evolution model). The reason we may attempt to answer this question in the case of the two systems under considerations is that Gaia BH1 and Gaia BH2 are not impacted by the orbit changes due to the mass transfer events after SN. Gaia BH1 has the orbital separation and eccentricity as it was shaped by SN and Gaia BH2 orbit is only modulated by weak tidal forces.
@@
Our work consists of two main parts: in Sec. 2 we consider IBE channel and follow the evolution of the binaries which we classify as Gaia BH1-like and Gaia BH2-like to find the probability of formation of such binaries from IBE, we then run the evolution of a two well-matching Gaia BH1 and Gaia BH2 binaries till the Hubble time is reached and finally we investigate the information that Gaia BH1 and Gaia BH2 may bring us about the natal kicks; in Sec. 3.1 we present and discuss the formation rates of Gaia BH-like binaries from young massive and open star cluster.
@@
Table 1: The parameters of Gaia BH1 and Gaia BH2: the orbital period $P_{\mathrm{orb}}$ in days, the eccentricity $e$, BH mass $M_{\mathrm{BH}}$ and companion star mass $M_{\star}$ in solar units, companion evolutionary stage and age in Gyrs, adopted from El-Badry et al. (2023a, b). The peculiar velocity at birth $v_{\mathrm{pec}}$ is taken from Table 3 in Zhao et al. (2023).
@@
$P_{\mathrm{orb}}$ [d] $185.59\pm 0.05$ $1276.7\pm 0.6$
$e$ $0.451\pm 0.005$ $0.5176\pm 0.0009$
$M_{\mathrm{BH}}$ [$\mathrm{M_{\odot}}$] $9.62\pm 0.18$ $8.94\pm 0.34$
@@ 2.1 Method
In our calculations the common envelope phase is parametrized by CE parameter $\alpha_{\mathrm{CE}}=\alpha_{\mathrm{ce}}\times\lambda$ where $\lambda$ is a parameter describing the donor central concentration (de Kool, 1989) and $\alpha_{\mathrm{ce}}$ is parametrizing the efficiency at which the orbital energy is transferred into the envelope (Webbink, 1984). We assume standard value $\alpha_{\mathrm{ce}}=1.0$ and for $\lambda$ we use the fits from Xu & Li (2010a); Xu & Li (2010b) and Wang et al. (2016).
@@ 2.2.2 The evolution of Gaia BH1- and Gaia BH2-like systems
The giant star loses its envelope becoming a Wolf-Rayet star on a close orbit ($a=140.3\,\mathrm{R_{\odot}}$) with MS star. In a subsequent evolution lasting $\sim 0.22$ Myr strong winds drive further mass loss from the primary star what widens the orbit to $a=164.4\,\mathrm{R_{\odot}}$ just prior the supernova (SN) explosion. The SN explodes $\sim 5.7$ Myr from the onset of the binary evolution. The primary star becomes $9.7\,\mathrm{M_{\odot}}$ BH and the effective kick (the composition of Blaauw and natal kicks) imparted to the system changes its orbital parameters to $a=318.8\,\mathrm{R_{\odot}}$ and $e=0.49$. Together with the MS primary of mass $1.2\,\mathrm{M_{\odot}}$ the system becomes Gaia BH1-like binary.
Our synthetic Gaia BH1 binary evolution proceeds on the nuclear timescale of the secondary which slowly evolves towards the end of hydrogen burning in its core. It takes $\sim 5$ Gyr for the star to leave the main sequence and become a red giant what marks the end of Gaia BH1 stage of its evolution. As the radius of red giant expands the tides act to circularize the orbit. Around $\sim 600$ Myr after the star became a red giant its radius reaches the radius of the star Roche lobe and the mass transfer onto BH begins. The system turns into low-mass X-ray binary (LMXB) of orbital period $P_{\mathrm{orb}}\sim 139.8$ days ($a=251.2\,\mathrm{R_{\odot}}$) and donor mass $M_{2}=1.2\,\mathrm{M_{\odot}}$. The mass transfer proceeds from less to more massive binary component, therefore, the orbit widens and so does the radius of the secondary star Roche lobe. When the Roche lobe extends to the point that the star fits inside it again the mass transfer ceases. It happens when the orbital separation is $a=891.7$ $\mathrm{R_{\odot}}$, $M_{\mathrm{BH}}\sim 10.2$ $\mathrm{M_{\odot}}$ and $M_{2}\sim 0.6$ $\mathrm{M_{\odot}}$. The LMXB phase lasts for $\sim 4$ Myr. From there the secondary star continues its nuclear evolution. Because the mass of the star is below the limit for helium ignition the red giant transforms into the helium white dwarf (WD) ($M_{\mathrm{WD}}=0.5$ $\mathrm{M_{\odot}}$ ) after shedding its diluted hydrogen envelope in stellar wind. At the Hubble time $t_{\mathrm{Hub}}=13.7$ Gyr the distance between the two compact objects in the system is $a=1244.3$ $\mathrm{R_{\odot}}$. For the rest of its life BH-helium WD binary slowly spirals in due to the gravitational radiation. The system will not merge within Hubble time.
@@
The system consisting of BH-RG forms $\sim 6.2$ Gyr after the onset of the binary evolution. While RG expands the tides act to tighten and circularize the orbit. When the orbital separation and the eccentricity decrease to $a=1175.3\,\mathrm{R_{\odot}}$ and $e=0.59$ respectively, they fall into the ranges which define the binary as Gaia BH2-like (see Table 2). This phase of the binary evolution lasts only $\sim 2$ Myr. In next $\sim 4$ Myr the eccentricity drops to zero and the orbital separation reaches its minimum, i.e. $a=757.7\,\mathrm{R_{\odot}}$. At this point the wind from the expanding RG comes into play: the orbit widens again, and BH accrets part of the matter lost by the evolved companion. By the time He is ignited in the core ($\sim 6.7$ Myr after orbit circularization) the star mass drops to $M_{2}=0.75\,\mathrm{M_{\odot}}$, BH mass slightly increases to $M_{\mathrm{BH}}=9.4\,\mathrm{M_{\odot}}$ and orbit widens to $a=1165.1\,\mathrm{R_{\odot}}$. The exhaustion of He in the core occurs after $\sim 112$ Myr and the companion becomes an early AGB star (E-AGB) and after next $\sim 5.7$ Myr it evolves to become the thermally-pulsing AGB star (TP-AGB). The AGB stage, when the star loses roughly $25\%$ of its mass and the orbit widens by almost a factor of two reaching $a=2216.9\,\mathrm{R_{\odot}}$, is followed by to the formation of carbon oxygen WD (CO WD) of mass $M_{\mathrm{WD}}=0.57\,\mathrm{M_{\odot}}$ orbiting BH of mass $M_{\mathrm{BH}}=9.5\,\mathrm{M_{\odot}}$ at a distance $a=2246.6\,\mathrm{R_{\odot}}$.
@@ 2.3.1 The formation rate of Gaia BH-like binaries
If we scale down IBE simulations to match the total simulated initial cluster mass of star clusters simulations $M_{\rm cl,tot}$$=3.6\times 10^{6}\,\mathrm{M_{\odot}}$ (see Sec. 3.2) we find one Gaia BH-like binary per simulation. This corresponds to the result of the dynamical interaction channel where also one Gaia BH0like binary ejected from the cluster was found in a given model cluster set. In conclusion, both channels give the same formation rate that is $\sim 2.7\times 10^{-7}\,\mathrm{M_{\odot}}^{-1}$.
2.3.2 The impact of natal kicks on Gaia BH-like binaries

### v2

# Kotko,I.,1 Banerjee,S.,2,3 Belczynski,K.,1 1Nicolaus Copernicus Astronomical Centre of the Polish Academy of Sciences, ul. Bartycka 18, 00-716 Warszawa, Poland

@@ Abstract
The two systems, namely, Gaia BH1 and Gaia BH2, that have been confirmed as dormant (i.e., no X-ray emission detected) black hole (BH) - low-mass star binaries in the latest Gaia mission data release (DR3) are intriguing in the context of their formation and evolution. Both systems consist of $\sim 9\,\mathrm{M_{\odot}}$ BH and $\sim 1\,\mathrm{M_{\odot}}$ star orbiting each other on a wide, eccentric orbit ($e\sim 0.5$). We argue that formation of such Gaia BH-like systems through the isolated binary evolution (IBE) channel, under the standard common envelope assumptions, and from dynamical interactions in young massive and open clusters are equally probable, and that the formation rate of such binaries is of the order of $10^{-7}\,\mathrm{M_{\odot}}^{-1}$ for both channels. We estimate that, according to our models, there are at most $\sim 900$ detectable Gaia BH-like binaries in the Milky Way thin disc. What plays an important role in formation of Gaia BH-like systems via the IBE channel is the mutual position of the natal kick velocity vector and the binary angular momentum vector. We find that natal kicks with a median magnitude of $\sim 40$ km/s are preferred for the formation of Gaia BH1-like binaries. Approximately $94\%$ of those binaries are formed with the BH spin misaligned to the orbital axis by less than $40^{\circ}$. Gaia BH2-like binaries form if the low velocity natal kick (of median magnitude $\sim 20$ km/s) is directed within $15^{\circ}$ about the orbital plane. In addition to natal kick, we also discuss the influence of tidal interaction and the adopted common envelope $\lambda_{\mathrm{ce}}$ parameter prescription on the evolution of Gaia BH-like binaries. We follow the subsequent evolution of the binaries, once formed as Gaia BH1 and Gaia BH2 systems, to investigate their connection with the low-mass X-ray binary population.
@@ 1 Introduction
The dormant black hole (BH) binaries are difficult to observe as they do not emit X-ray radiation and can be discovered only by astrometric and/or spectroscopic analysis. Although such systems were predicted to exist already in the 1960s (Guseinov & Zel’dovich, 1966), the new generation of the astronomical instruments and new data analysis methods have allowed to identify the first candidates for such objects only recently. To our knowledge, the first strong candidate for a binary system with a heavy companion unseen in electromagnetic spectrum was found in the globular cluster NGC 3201 Giesers et al. (2018). Within the last two years, another four systems have been discovered: (i) binary HD 130298 where the minimum mass of an unseen companion has been estimated to be $7.7\pm 1.5$ $\mathrm{M_{\odot}}$ (Mahy et al., 2022), (ii) VFTS 243 found in Large Magellanic Cloud which is a binary of a massive star and most probably a black hole companion with mass $M=10.1\pm 2.0$ $\mathrm{M_{\odot}}$ (Shenar et al., 2022), and (iii) two systems which were identified in the most recent Gaia catalogue DR3 for which there is the strong evidence that they host a non-accreting BH and a low-mass companion star (Chakrabarti et al., 2023; El-Badry et al., 2023a, b; Tanikawa et al., 2023b). In our studies, we focus only on the last two systems and, following El-Badry et al. (2023b), we refer to those binaries as Gaia BH1 and Gaia BH2.
@@
Once Gaia BH1 and Gaia BH2 had been detected, questions and speculations arose about the origin of those systems. Several studies have already been made which point towards dynamical interactions in young stellar clusters as the most probable scenario to create a BH-low-mass star system having a moderately wide and eccentric orbit (El-Badry et al., 2023a; Tanikawa et al., 2023a; Rastello et al., 2023; Niccolò Di Carlo et al., 2023). However, as we show in this work, the isolated binary evolution (IBE) channel may contribute to the total population of dormant BH binaries at a rate that is comparable to that for the dynamical-interaction channel. The important factors that impact the number of the binaries of interest born through the IBE channel is the magnitude and the geometry of the natal kick imparted to the newborn compact object after the supernova (SN) explosion. One can infer about the natal kick from the peculiar velocity [^1] The peculiar velocity corresponds to systemic velocity in the model, as is explained in Sec. 2.1. and the orbital parameters of the binary system.
@@
There are two components that change the systemic velocity of a binary after an SN event in the model: (i) the Blaauw kick that impacts the velocity of the binary centre of gravity due to symmetric mass ejection during the SN (Blaauw, 1961), (ii) the natal kick that is received by the compact object due to asymmetries of mass ejecta and/or neutrino emission during the SN (Chugai, 1984; Janka, 2013, 2017, e.g.). The observations of proper motions of pulsars and X-ray binaries give constrains on the natal kicks received by neutron stars (NSs) or BHs upon their formation. The Maxwellian distribution with dispersion velocity $\sigma=265$ km/s, as derived by Hobbs et al. (2005) from observations of Galactic pulsars, is widely used in the binary population synthesis simulations. This distribution is applied also to BHs but with additional assumption that the kick velocity is reduced by the amount of matter that falls back, being pulled by the gravitational field of the BH forming during the SN. However, recent observations indicate that this approach may underestimate the number of lower velocity systems and overestimate the number of the systems in higher velocity range and that it should be revisited (Atri et al., 2019; O’Doherty et al., 2023; Zhao et al., 2023). The magnitude of the natal kicks is only one part of the puzzle, the other is the kick velocity vector orientation. In the codes based on BSE (Hurley et al., 2000), the direction of a natal kick is drawn from an isotropic spherical distribution. The question is whether the specific direction of the natal kick, that a BH receives in an SN explosion, could lead to the formation of Gaia BH1 and Gaia BH2 (under the standard assumptions of isolated binary evolution model). The reason we may attempt to answer this question for the two systems of interest is that Gaia BH1 and Gaia BH2 are not impacted by the orbit changes due to the mass transfer events after the SN. Gaia BH1 has the orbital separation and eccentricity as it was shaped by SN and Gaia BH2 orbit is only modulated by weak tidal forces.
@@
Our work consists of two main parts: in Sec. 2, we consider the IBE channel and follow the evolution of the binaries which we classify as Gaia BH1-like and Gaia BH2-like and find the probability of formation of such binaries from the IBE. We then run the evolution of two well-matching Gaia BH1 and Gaia BH2 binaries until the Hubble time and investigate what information Gaia BH1 and Gaia BH2 may bring us about the natal kicks. In Sec. 3, we present and discuss the formation rates of Gaia BH-like binaries from young massive and open star clusters.
@@
Table 1: The parameters of Gaia BH1 and Gaia BH2: the orbital period $P_{\mathrm{orb}}$ in days, the eccentricity $e$, BH mass $M_{\mathrm{BH}}$ and companion star mass $M_{\star}$ in solar units, companion evolutionary stage and age in Gyrs, adopted from El-Badry et al. (2023a); Chakrabarti et al. (2023); El-Badry et al. (2023b). The peculiar velocity at birth $v_{\mathrm{pec}}$ is taken from Table 3 in Zhao et al. (2023).
@@
$P_{\mathrm{orb}}$ [d] $185.52\pm 0.08$ $1276.7\pm 0.6$
$e$ $0.439^{0.004}_{-0.003}$ $0.5176\pm 0.0009$
$M_{\mathrm{BH}}$ [$\mathrm{M_{\odot}}$] $9.326^{0.216}_{-0.209}$ $8.94\pm 0.34$
@@ 2.1 Method
In our calculations the common envelope phase is parameterized by CE parameter $\alpha_{\mathrm{CE}}=\alpha_{\mathrm{ce}}\times\lambda_{\mathrm{ce}}$ where $\lambda{\mathrm{ce}}$ is a parameter describing the donor central concentration (de Kool, 1989) and $\alpha_{\mathrm{ce}}$ parameterizes the efficiency at which the orbital energy is transferred into the envelope (Webbink, 1984). We assume the standard value $\alpha_{\mathrm{ce}}=1.0$ and for $\lambda_{\mathrm{ce}}$ we use the fits from Xu & Li (2010a); Xu & Li (2010b) and Wang et al. (2016) (see Sec.2.3.3 for discussion).
@@ 2.2.2 The evolution of Gaia BH1- and Gaia BH2-like systems
The giant star loses its envelope becoming a Wolf-Rayet star on a close orbit ($a\sim 140\,\mathrm{R_{\odot}}$) with MS star. In the subsequent evolution that lasts $\sim 0.22$ Myr, the strong winds drive further mass loss from the primary star and the orbit expands to $a\sim 164\,\mathrm{R_{\odot}}$ just prior the supernova (SN) explosion. The SN explodes $\sim 5.7$ Myr from the onset of the binary evolution. The primary star becomes $9.7\,\mathrm{M_{\odot}}$ BH and the effective kick (the composition of Blaauw and natal kicks) imparted to the system increases both its orbital separation and the eccentricity to $a\sim 319\,\mathrm{R_{\odot}}$ and $e=0.49$ leading to the formation of Gaia BH1-like binary.
Our synthetic Gaia BH1 binary evolution proceeds on the nuclear timescale of the secondary which slowly evolves towards the end of hydrogen burning in its core. It takes $\sim 5$ Gyr for the star to leave the main sequence and become a red giant what marks the end of Gaia BH1 stage of the binary evolution. As the radius of the red giant expands the tides act to circularize the orbit. About $\sim 600$ Myr later the red giant radius reaches the radius of the star Roche lobe and the mass transfer onto BH begins. The system turns into low-mass X-ray binary (LMXB) with the orbital period $P_{\mathrm{orb}}\sim 140$ days ($a\sim 251\,\mathrm{R_{\odot}}$) and donor mass $M_{2}=1.2\,\mathrm{M_{\odot}}$. The mass transfer proceeds from less to more massive binary component, therefore, the orbit widens and so does the radius of the secondary star Roche lobe. When the Roche lobe extends to the point that the star fits inside it again, the mass transfer ceases. It happens when the orbital separation is $a\sim 892$ $\mathrm{R_{\odot}}$, $M_{\mathrm{BH}}\sim 10.2\,\mathrm{M_{\odot}}$ and $M_{2}\sim 0.6\,\mathrm{M_{\odot}}$. The LMXB phase lasts for $\sim 4$ Myr. From there, the secondary star continues its nuclear evolution until it transforms into a carbon-oxygen white dwarf (WD) of mass $M_{\mathrm{WD}}=0.5\,\mathrm{M_{\odot}}$. At the Hubble time $t_{\mathrm{Hub}}=13.7$ Gyr, the distance between the two compact objects in the system is $a\sim 1244\,\mathrm{R_{\odot}}$. For the rest of its life BH–WD binary slowly spirals in due to the gravitational radiation. The system will not merge within Hubble time.
@@
The system consisting of BH–RG forms $\sim 6.2$ Gyr after the onset of the binary evolution. While RG expands the tides act to tighten and circularize the orbit. When the orbital separation and the eccentricity decrease to $a\sim 1175\,\mathrm{R_{\odot}}$ and $e=0.59$ respectively, they fall into the ranges which define the binary as Gaia BH2-like (see Table 2). This phase of the binary evolution lasts only for $\sim 2$ Myr. In next $\sim 4$ Myr the eccentricity drops to zero and the orbital separation reaches its minimum, i.e. $a\sim 758\,\mathrm{R_{\odot}}$. At this point the wind from the expanding RG comes into play: the orbit widens again, and BH accrets part of the matter lost by the evolved companion. By the time He is ignited in the core ($\sim 6.7$ Myr after orbit circularization) the star mass drops to $M_{2}=0.75\,\mathrm{M_{\odot}}$, BH mass slightly increases to $M_{\mathrm{BH}}=9.4\,\mathrm{M_{\odot}}$ and orbit widens to $a\sim 1165\,\mathrm{R_{\odot}}$. The exhaustion of He in the core occurs after $\sim 112$ Myr when the companion becomes an early AGB star (E-AGB) and after next $\sim 5.7$ Myr it evolves to become the thermally-pulsing AGB star (TP-AGB). The AGB stage, when the star loses roughly $25\%$ of its mass and the orbit widens by almost a factor of two reaching $a\sim 2217\,\mathrm{R_{\odot}}$, is followed by the formation of carbon oxygen WD (CO WD) of mass $M_{\mathrm{WD}}=0.57\,\mathrm{M_{\odot}}$ orbiting BH of mass $M_{\mathrm{BH}}=9.5\,\mathrm{M_{\odot}}$ at a distance $a\sim 2247\,\mathrm{R_{\odot}}$.
@@ 2.3.1 The formation rate of Gaia BH-like binaries
If we scale down IBE simulations to match the total simulated initial cluster mass of star clusters simulations $M_{\rm cl,tot}$$=3.6\times 10^{6}\,\mathrm{M_{\odot}}$ (see Sec. 3.2) we find one Gaia BH-like binary per simulation. This corresponds to the result of the dynamical interaction channel where also one Gaia BH-like binary ejected from the cluster was found in a given model cluster set. In conclusion, both channels give the same formation rate that is $\sim 2.7\times 10^{-7}\,\mathrm{M_{\odot}}^{-1}$.
2.3.2 The estimated number of Gaia BH-like binaries
A considerable amount of work has so far been done to estimate the number of dormant BHs that can be present in the Milky Way within Gaia detection capability (e.g. Breivik et al. (2017); Mashian & Loeb (2017); Yamaguchi et al. (2018); Wiktorowicz et al. (2020); Shikauchi et al. (2022); Chawla et al. (2022); Shikauchi et al. (2023)). Typically, the resulting number of binaries of interest varies from few tens to few hundreds (Yamaguchi et al., 2018; Wiktorowicz et al., 2020; Chawla et al., 2022, e.g.). However, the number can reach as high as $10^{3}-\sim 10^{5}$ (Breivik et al., 2017; Mashian & Loeb, 2017, e.g.) depending on the physics adopted in the evolutionary models, on the accuracy with which the Milky Way is modelled (where for example the realistic stellar distribution (Wiktorowicz et al., 2020; Chawla et al., 2022, e.g.)), and/or how the position-dependent reddening (Chawla et al., 2022) is taken into account.
Although the main focus of our paper is the comparison of the formation rates of Gaia BH-like binaries for the isolated-binary and the dynamical formation channels, we complement this work with an approximate estimation of the expected number of Gaia BH-like binaries in the Milky Way within the scope of Gaia detectability, according to the two models. We limit our considerations to the Galactic thin disc population, since our simulations of the isolated binary evolution are done for a metallicity that is close to solar.
As an intrinsic synthetic population of Gaia BH-like binaries, we choose that from the model “V2 low-mass comp” (see 2.2.1), where we assume that the natal kicks distribution is a double peaked Maxwellian and require the mass of the companion to be smaller than $3.0\,\mathrm{M_{\odot}}$ for the binary to be classified as Gaia BH-like. We calibrate the mass of the simulations to match the stellar mass of the Galactic thin disc, where, for the latter, we adopt $M_{\mathrm{d,thin}}=2.15\times 10^{10}\,\mathrm{M_{\odot}}$ following Breivik et al. (2017) who assumed a constant star formation rate of $2.15\,\mathrm{M_{\odot}}\mathrm{yr}^{-1}$ over 10 Gyr. We draw the present-day population of BH–low-mass star binaries from this intrinsic binary population, taking into account the estimated age of Gaia BH1 ($\sim 8$ Gyr) and the estimated duration of the red giant phase of the companion star in Gaia BH2 ($\sim 100$ Myr).
In the next step, the resulting binary population is spatially distributed at random, following the thin disc mass distribution described by the function from Yu & Jeffery (2015), i.e.,
$\rho(R,z)=\frac{M_{\mathrm{d,thin}}}{4\pi h_{R}\,h_{z}}e^{R/h_{R}}\,sech^{2}(-z/h_{z}),$ (1)
where $h_{R}=2.5$ kpc is the scale lenght of the disc, $h_{z}=0.352$ kpc is its scale height, and $R$ and $z$ are the cylindrical coordinates of the axisymmetric disc.
To calculate the apparent magnitude $m_{V}$ of all considered present-day Gaia BH-like binaries, we first calculate the absolute magnitude and the bolometric correction (Torres, 2010) for each binary system in the sample using the luminosity, $L_{2}$, and the effective temperature, $T_{\mathrm{eff,2}}$, of the companion star. For the interstellar extinction in the $V$-band, $A_{V}$, we follow the approach of Yamaguchi et al. (2018) who assumed that $A_{V}(d)=d$. Here, $d$ in the distance from Earth in kpc that we calculate from the binary’s position in the thin disc. We set a lower limit of $m_{V}<20$ mag for the apparent magnitude of the binary, if it is to be detected by Gaia.
The astrometric signature, $\alpha$, is calculated following Breivik et al. (2017):
$\alpha=\biggl{(}\frac{a_{\mathrm{proj}}}{\mathrm{au}}\biggr{)}\biggl{(}\frac{d}{\mathrm{pc}}\biggr{)}^{-1}\,\,\,\mathrm{arcsec}$ (2)
where $d$ is the distance to the binary and $a_{\mathrm{proj}}$ is the projected semi-major axis of the luminous companion.
We calculate $a_{\mathrm{proj}}$ for each system, using the exact position of the star on its orbit around the compact object and the orbital elements $i$, $\Omega$ and $\omega$ at a randomly chosen time elapsed from the star’s periastron passage. We consider the binary as detectable if $\alpha>3\sigma_{\mathrm{G}}$, where $\sigma_{\mathrm{G}}$ is Gaia’s astrometric precision calculated following Gaia Collaboration et al. (2016) (see Eq.4 therein).
The final estimated number of BH–low-mass star binaries expected in the Galactic thin disc from the model of isolated binary evolution, assuming the most optimistic classification criteria, is of the order of $1.4\times 10^{3}$ and, among them, $\sim 900$ binaries are within Gaia’s detectability limit. In Fig. 5, we present the distributions of the orbital periods and eccentricities of those binaries. It can be seen that binaries with orbital period shorter than $100.0$ days ($\sim 66\%$ of all binaries) and eccentricity larger than $0.5$ ($\sim 68\%$ of all binaries) dominate the distributions. About half of the binaries ($\sim 44\%$) have both $P_{\rm orb}$ and $e$ in these ranges.
Figure 5: The distribution of the orbital periods $P_{\mathrm{orb}}$ (top) and eccentricity $e$ (bottom) of the predicted population of Gaia BH-like binaries evolved through IBE detectable by Gaia during 5 yr observational period.
2.3.3 The evolution Gaia BH-like binaries
In the following paragraphs we focus on the three phases of the binary evolution that we find important to discuss. The thorough studies of the impact of the choice of SN model (rapid vs delayed, with or without fallback), or of changing the value of the common envelope efficiency parameter $\alpha_{\mathrm{ce}}$ on the formation of detached BH–star binaries can be found for example in Chawla et al. (2022) and Shikauchi et al. (2023). Although our results differ from conclusions of Shikauchi et al. (2023) on $\alpha_{\mathrm{ce}}$ (see Sec. 2.3.3) it is instructive to see the influence of $\alpha_{\mathrm{ce}}>1$ on the dormant BH populations.
The evolution Gaia BH-like binaries: the impact of the natal kicks.